OpenStudy (anonymous):

just to make sure... (e^(-2t))^2 = e^(4t) right?

OpenStudy (anonymous):

no, it is e^(-4t)

OpenStudy (agreene):

-2^2 = -4 so agdgdgdgwngo is correct.

OpenStudy (anonymous):

er yeah negative

OpenStudy (agreene):

-2^2 = -4 (-2)^2 = 4 order of operations ;)

OpenStudy (anonymous):

oh right :-P

OpenStudy (anonymous):

hmm wait a minute, I thought unary minus has precedence over the binary exponentiation

OpenStudy (anonymous):

so -2^2 = (-2)^2 = 4

OpenStudy (anonymous):

but my calculator says exponentiation has higher precedence... so :-(

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=-2%5E2

OpenStudy (agreene):

I believe, exponentiation has almost the highest order. Also, the parenthetical statement can be said as such: -2^2 = -1(2^2) = -1(4) = -4 Whereas: (-2)^2 = (-2)*(-2) = 4

OpenStudy (jamesj):

For the record, \[(a^x)^y = a^{xy} \ \ \text{ NOT } \ \ a^{x^y}\]

OpenStudy (jamesj):

Hence \[(e^{-2x})^2 = e^{-4x}\]